a ceiling fan rotating at its full speed of 1200rpm is swithched off at t =0 .it cums at rest at t=5 s assuming uniform motion ,the no. of revolution made by it b4 cuming to rest ??
1
Expert's answer
2012-12-10T12:00:06-0500
Equation of motion is
f = wt-et^2/2
where f is number of revolutions, t is time, w is initialrotation speed and e is angular deceleration.
it is obvious that e = w/(1/12 of minute) = 14400minute^(-2)

Hence
f = 1200t-14400t^2/2, t=1/12 of minute = 5 sec.
f = 100-50 = 50 revolutions will be done.